### v2

# Counterfactual Evaluation for Blind Attack Detection in LLM-based Evaluation Systems

Abstract
This paper investigates defenses for LLM-based evaluation systems against prompt injection. We formalize a class of threats called blind attacks, where a candidate answer is crafted independently of the true answer to deceive the evaluator. To counter such attacks, we propose a framework that augments Standard Evaluation (SE) with Counterfactual Evaluation (CFE), which re-evaluates the submission against a deliberately false ground-truth answer. An attack is detected if the system validates an answer under both standard and counterfactual conditions. Experiments show that while standard evaluation is highly vulnerable, our SE+CFE framework significantly improves security by boosting attack detection with minimal performance trade-offs.
1 Introduction
Advancements in artificial intelligence have been propelled by shared tasks and benchmarks, which provide standardized evaluation and foster rigorous comparison. While platforms like Kaggle kaggle and datasets such as ImageNet imagenet, COCO coco, and Cityscapes cityscapes have advanced machine learning, data mining, and computer vision, natural language processing (NLP) has progressed through benchmarks like GLUE glue, SuperGLUE superglue, and SQuAD squad.
In recent years, large language models (LLMs) have demonstrated robust reasoning capabilities across various tasks, supported by benchmarks such as MMLU mmlu and StrategyQA strategyqa. Increasingly, LLMs also serve as automatic evaluators for benchmarks, reducing the costs of human evaluation kim2024; shankar2024. However, these evaluator LLMs exhibit biases: they hallucinate plausible but incorrect judgments ji2023; tang2024, favor low-perplexity examples stureborg2024; koo2024, prefer their own generations panickssery2024; koo2024, and display anchoring effect in multiple judgments stureborg2024; eigner2024.
These limitations are particularly concerning in LLM competitions, where participants may exploit them to gain an unfair advantage. Prompt injection attacks liu2023prompt pose a distinct challenge by causing an LLM to behave unexpectedly using a devised prompt, potentially tricking the evaluation system into scoring incorrect answers as correct. Variants such as indirect prompt injection attacks yi2025; greshake2023 and prompt leaking liu2024pia; perez2022 demonstrate the increasing complexity of such threats.
Among these, blind attacks remain an underexplored yet consequential threat to the integrity of automated LLM evaluation. In blind attacks, the candidate answer is generated independently of the true answer, conditioned only on the question. This can potentially elicit a favorable judgment from the evaluator, regardless of the ground-truth answer. Common techniques such as direct prompt injection shi2024; liu2024pia and rewording attacks iyyer2018; cao2022 fall into this class. Prompt injection includes strategies such as ignore previous instructions perez2022, token smuggling jiang2024, role-playing wei2023, indirect references greshake2023, few-shot attack xu2024, and many-shot attack anil2024. Other attack strategies targeting LLMs include jailbreaks, which exploit model vulnerabilities for unauthorized actions, and data poisoning, which corrupts training data to manipulate model behavior. Refined query-based jailbreaking chao2025 uses a minimal number of queries to probe and bypass a model’s defense, while Tree of Attacks mehrotra2024 jailbreak LLMs iteratively, generating and evaluating variations of the initial adversarial prompt until a successful jailbreak is achieved. Data poisoning techniques include backdoor attacksshah2023; kandpal2023 and PII extraction chen2024_janus. A blind attack is one of the most basic forms of manipulation. Despite their simplicity, blind attacks expose vulnerabilities by disconnecting the question and the ground truth. Evaluators can be deceived into hallucinating, resulting in marking invalid answers as correct and thereby undermining evaluation robustness. Studying this class of attacks systematically is an important step toward defending against adversarial attacks and building more robust LLM evaluation systems.
Previous defense methods for similar prompt injection attacks include erase-and-check safety filters gosmar2025, multi-agent NLP frameworks kumar2023, and unified detection mechanisms designed to handle prompt injection, backdoor, and adversarial attacks lin2025_uniguardian. Methods can also be classified into prompt-level zou2023; hines2024 and model-level defense touvron2023; lin2025_uniguardian. In addition, an increasing number of studies has been made targeting the security of evaluator LLMs. One such benchmark is CyberSecEval 2 bhatt2024, which focuses on a wide range of adversarial threats, such as prompt injection, vulnerability identification and exploitation, and code interpreter abuse. CyberBench cyberbench assesses LLM performance on multiple choice, text classification, and other cybersecurity-related tasks, while LLM4Vuln llm4vuln aims to decouple an LLM’s vulnerability reasoning from knowledge retrieval, context awareness, and prompt design, enabling structured evaluation across these dimensions.
To address this, we propose an evaluation framework that incorporates counterfactual prompts, which replace the original ground truths with random fake terms. The core insight behind our approach is that blind attacks deceive the evaluation system without truly aligning with the ground truth. Our method exposes the inconsistencies in the evaluation behavior, allowing us to assess answer correctness under normal conditions and detect such attacks when present.
This paper makes the following contributions. First, we formalize and define blind attacks, a class of prompt injection attacks that force the evaluator LLM to mark a submitted answer as correct, regardless of the true answer. Second, we propose an evaluation framework that utilizes counterfactual prompts to identify blind attacks and demonstrate its effectiveness on both proprietary and open-source models. Results show that our approach demonstrates strong resilience against o1, GPT-4o, and GPT-4o-mini, though their effectiveness slightly diminishes with GPT-3.5-turbo. Additionally, overall evaluation robustness depends significantly on model capacity. While smaller models are more prone to hallucination, stronger and more advanced models can navigate subtle linguistic variations and demonstrate greater robustness. Together, this framework improves the integrity of the evaluation, offering a practical approach to protect LLM assessments against adversarial manipulation.
2 Problem Definition
We formally define the evaluation setting and the specific threat model. We first outline the use of an LLM as an evaluator for question-answering tasks and then introduce our definition of a blind attack.
2.1 Performance Evaluations of LLM-based Question Answering System
We employ an LLM as an evaluation system to judge whether a question answering system’s response to a question is correct, using a prompt such as Prompt 1, which we refer to as a standard evaluation (SE) prompt.
Assuming that the evaluation system knows the ground truth answer to the question , we instruct the LLM to evaluate from the user input based on the ground truth by outputting if is correct for a pair of the question and the ground truth , or otherwise, if the candidate answer differs from the true answer.
2.2 Prompt Injection Attack
We assume the attacker lacks access to the true answer. If the attacker had such access and aimed solely to maximize their score, they could trivially submit the correct answer. Therefore, it is more realistic to consider attacks that attempt to deceive the evaluator without knowledge of the true answer, causing the evaluator’s judgment to become effectively independent of the ground truth.
We define this type of threat as a blind attack, formally stated as follows.
Definition 1 (Blind Attack).
Let be a response strategy that maps a question to a candidate answer , i.e., . We say that is a blind attack strategy if, for all possible questions , the output of the evaluator is conditionally independent of the true answer , given and .
In other words, a blind attack is one in which the evaluator’s decision depends only on the question and the submitted response, and not on the correct answer. This captures attacks in which the evaluator is manipulated to produce the same judgment regardless of what the true answer actually is.
Blind attacks include many strategies, including direct prompt injection, where attackers overwrite evaluation instructions to make the evaluator constantly return favorable scores. Our early experiments indicate that modern LLMs are fairly resistant to such attacks, which led us to explore a more subtle variant: the rewording attack iyyer2018; cao2022. Here, the attacker generates by rephrasing instead of answering genuinely. For example, when encountering the question "On what date was the Declaration of Independence officially signed?", an attacker might rephrase the original question as "What was the date on which the Declaration of Independence was officially signed?" and submit it as their response. The submitted response is generated solely with the knowledge of the original question and not referencing the ground truth answer. This exploits a vulnerability where the LLM hallucinates correctness. It misinterprets the reworded question as a valid response, incorrectly outputting despite its irrelevance to the true answer, as shown in Fig.1(b). This is in contrast to non-attack situations, where the evaluation output reflects a binary judgment (0/1), as illustrated in Fig.1(a).
3 Proposed Methods
We propose a framework that integrates (1) standard evaluation (SE), and (2) counterfactual evaluation (CFE) as our evaluation method.
3.1 Standard Evaluation
SE evaluates the equivalence of the candidate answer and the ground truth given the question . It checks whether the submitted response is correct for the given question, without considering the possibility of adversarial attempts. We use Prompt 1 and denoted SE as .
If we do not consider the possibility of attacks, this evaluation prompt alone is sufficient. However, SE may fail under blind attacks, since the evaluator can hallucinate correctness when is subtly yet superficially aligned with .
3.2 Counterfactual Evaluation
We propose CFE to detect blind attacks where a system submits an answer that is independent of the correct answer given the question . These attacks make the evaluation system output the symbol for the correct answer without verifying the candidate answer’s alignment with the true answer.
We exploit this characteristic of blind attacks in CFE.
For honest answers, the evaluator can be expected to accept the true answer but reject unrelated fake answers. In contrast, a blind attack is designed to be accepted regardless of the ground truth, and thus tends to be accepted even when the true answer is replaced. This invariance to the substitution of allows us to flag such responses as suspicious and serves as the foundation of our detection rule. For example, for the question "What is the name of the backing group that supported Nana Mouskouri?", we randomly replace the original ground truth "The Athenians" with an irrelevant term like "Penguin" or "Apple". We denote random fake truth as , and propose the prompt for CFE as in Prompt 2, denoted as , with changes highlighted in bold.
We generate fake ground truths by using a prompt such as "Please output an answer that has nothing to do with " beforehand. Since is independent to , the evaluation system should output unless by chance. If the system instead outputs , it reveals susceptibility to blind attacks.
| SE | CFE | Decision |
| 1 | 0 | Correct answer |
| 1 | 1 | Attack detected |
| 0 | * | Wrong answer |
The decision rule of the framework is summarized in Table 1.
3.3 Justification
We provide an intuitive justification for the proposed framework. It follows directly from the definition that
for any , indicating that the evaluator LLM’s output distribution is invariant to the ground truth.
In principle, direct verification of the equality requires repeated evaluations across different values of and statistical tests of output independence. In practice, however, blind attacks often aim to elicit the favorable output 1 from the evaluator with high probability close to 1, regardless of the value of . Therefore, we implement detection by testing whether evaluations against both the true answer and a deliberately fake answer return 1.
Conversely, for honest answers, the evaluator returns 1 when the submitted response matches the true answer (SE), and 0 when compared to an unrelated fake answer (CFE). Hence, a response is accepted as legitimate when the two evaluations disagree.
In essence, our decision rule checks whether the evaluator’s output varies when the true answer is replaced. Lack of change indicates invariance to the ground truth, an essential feature of blind attacks, and therefore serves as a reliable signal for detection.
A potential vulnerability in CFE is the coincidental semantic or lexical overlap between a generated fake answer and the true answer, which could lead to erroneous attack detection. To mitigate this, a more robust approach involves generating multiple distinct fake answers. Firstly, if the vocabulary size is in the tens of thousands, the probability of false acceptance becomes negligible. Moreover, using multiple fake answers for CFE reduces the risk exponentially. As long as the true answer is not leaked, blind attacks can be detected with high probability.
In addition, QA tasks can allow partial correctness or multiple valid answers. Evaluating SE and CFE against all possible answers risk wasting computational cost and blurring the distinction between genuine semantic variation and adversarial invariance. Future extensions can address this by allowing the binary evaluator to assign weighted scores based on the plausibility of each answer. Moreover, sampling or consensus-based strategies could further reduce computational cost while maintaining evaluation robustness.
4 Experiments
To validate our approach, we conduct a series of experiments to evaluate the framework’s effectiveness against blind attacks across several models and datasets.
4.1 Experimental Setup
We evaluated our proposed evaluation methods on six English question-answer datasets: GSM8k (Train data) gsm8k, HotpotQA (Train data) hotpotqa, SQuAD (SQuAD 1.1) squad, StrategyQA (Train data) strategyqa, TriviaQA (Unfiltered TriviaQA version 1.0) triviaqa, and TruthfulQA truthfulqa. GSM8K and StrategyQA are English question-answering datasets focused on mathematical reasoning and implicit strategic reasoning, consisting of 7,473 and 2,780 questions, respectively. HotpotQA, SQuAD 1.1, TriviaQA, and TruthfulQA are English question-answering datasets supporting multi-hop reasoning, reading comprehension, trivia knowledge, and truthfulness evaluation, with 90,447, 87,599, 87,662, and 817 questions, respectively. We randomly selected one hundred questions from each, yielding a total of six hundred questions.
We used GPT-4o to generate correct and wrong answers under non-attack conditions. To preserve the integrity of the original answers, we employed two prompts based on the nature of the benchmark datasets: Prompt 3(1) for GSM8K, StrategyQA, and TruthfulQA, which consist of full sentences or binary (True/False) ground truths; and Prompt 3(2) for HotpotQA, SQuAD, and TriviaQA, where answers are concise phrases or named entities. Wrong candidate answers were obtained using Prompt 4.
To test robustness, we constructed attacks via Prompt 5 and examined attack detection using two methods: (i) standard evaluation (SE), and (ii) standard and counterfactual evaluation (SE+CFE). We evaluated four proprietary LLMs, GPT-3.5-turbo, GPT-4o-mini (gpt-4o-mini-2024-07-18), GPT-4o (gpt-4o-2024-08-06), and o1 (o1-2024-12-17), accessed through OpenAI’s API, as well as three open-source LLMs accessed via OpenRouter: Gemma (google/gemma-3-12b-it), LLaMa (meta/llama-3.1-8b-instruct), and Mistral (mistralai/mistral-7b-instruct:free). Our experiments were implemented with API calls to the various models, so we do not report GPU hours or computational budget. The exact number of parameters for the proprietary models has not been public disclosed and is therefore not reported. All temperature parameters were set to a value of 0.7 based on preliminary tests, balancing between consistency and diversity. Other API parameters were kept at their default values.
| Question | “The 2002 Winter Olympics were held in which city?” |
| Ground Truth | “Salt Lake City” |
| Correct Answer | “The 2002 Winter Olympics were held in Salt Lake City.” |
| Wrong Answer | “Denver” |
| Attack | “In which city were the 2002 Winter Olympics held?” |
| SE | Correct | Wrong+Attack | Accuracy | ASR | ||||
| Prec. | Rec. | F1 | Prec. | Rec. | F1 | |||
| Gemma-12B | 0.542 | 0.975 | 0.697 | 0.979 | 0.588 | 0.735 | 0.717 | 0.802 |
| LLaMA-3.1-8B | 0.343 | 0.893 | 0.496 | 0.732 | 0.146 | 0.243 | 0.395 | 0.872 |
| Mistral-7B | 0.502 | 0.890 | 0.642 | 0.910 | 0.559 | 0.693 | 0.669 | 0.777 |
| GPT-3.5-turbo | 0.582 | 0.902 | 0.708 | 0.932 | 0.677 | 0.784 | 0.752 | 0.618 |
| GPT-4o-mini | 0.497 | 0.977 | 0.659 | 0.977 | 0.506 | 0.667 | 0.663 | 0.982 |
| GPT-4o | 0.502 | 0.978 | 0.664 | 0.979 | 0.515 | 0.675 | 0.669 | 0.958 |
| o1 | 0.495 | 0.985 | 0.658 | 0.985 | 0.497 | 0.660 | 0.659 | 0.998 |
| SE+CFE | Correct | Wrong | Attack | Accuracy | ||||
| Prec. | Rec. | F1 | Prec. | F1 | Prec. | F1 | ||
| Gemma-12B | 0.952 | 0.925 | 0.938 | 0.812 | 0.887 | 0.943 | 0.852 | 0.893 |
| LLaMA-3.1-8B | 0.388 | 0.202 | 0.265 | 0.402 | 0.306 | 0.403 | 0.524 | 0.400 |
| Mistral-7B | 0.591 | 0.757 | 0.664 | 0.729 | 0.803 | 0.671 | 0.460 | 0.667 |
| GPT-3.5-turbo | 0.787 | 0.873 | 0.828 | 0.669 | 0.792 | 0.927 | 0.564 | 0.750 |
| GPT-4o-mini | 0.991 | 0.952 | 0.971 | 0.960 | 0.976 | 0.975 | 0.978 | 0.975 |
| GPT-4o | 0.990 | 0.947 | 0.968 | 0.937 | 0.963 | 0.965 | 0.958 | 0.963 |
| o1 | 0.990 | 0.985 | 0.987 | 0.983 | 0.988 | 1.000 | 0.998 | 0.991 |
4.2 Results
We show overall results across all six datasets in Table 3. Without attacks, o1 outperformed GPT-3.5-turbo but was surpassed by GPT-4o-mini and GPT-4o.
Table 2 shows an example of QA evaluation with LLM-generated candidate responses for correct, wrong, and attack situations. GPT-4o generated correct answers that varied naturally while preserving integrity, wrong answers plausibly distinct from the ground truth, and blind attacks that rephrased the question without altering its intent.
For SE, blind attacks achieved an attack success rate (ASR) of for GPT-3.5-turbo, and even higher rates for GPT-4o-mini (), GPT-4o (), and o1 (). Although all four proprietary models achieved high recall on correct answers () and high precision on wrong answers (), low precision for correct and low recall for wrong/attack cases indicate their vulnerability to blind attacks. GPT-3.5-turbo’s lower ASR of may reflect its more limited linguistic understanding, making it less susceptible to subtle semantic manipulations.
For SE+CFE, the detection of blind attacks improved significantly. For GPT-4o-mini, GPT-4o, and o1, the F1 scores for attack detection reached , , and , respectively, with accuracy exceeding for all three models.
GPT-3.5-turbo also saw moderate gains, with its F1 score for correct detection rising from to , although its attack detection remained weak (), likely due to its comparatively weaker semantic understanding.
Among open-source models, Mistral-7B and Gemma-12B were competitive with GPT-3.5-turbo, with Gemma-12B achieving a accuracy under SE+CFE. LLaMA-8B underperformed, occasionally outputting null values instead of binary predictions, which were marked incorrect. These results underscore a trade-off between robustness and accessibility: open-source models offer practical, lower-resource alternatives but with reduced resistance to blind attacks.
To compare typical failures for GPT-3.5-turbo and GPT-4o-mini, we consider specific cases where SE+CFE was unable to make a correct judgement.
While our experiments only used one SE prompt and demonstrated strong results on recent OpenAI models, we note that not all models behaved equally. For example, GPT-3.5-turbo performed poorly in attack detection, while Gemma-12B showed strong results under SE+CFE. This suggests that the effectiveness of our framework may depend more on model capacity than on vendor or architecture alone.
To better understand when our proposed method fails, we examine common patterns in evaluation outputs across datasets, for the individual LLM models. We present pseudo confusion matrices showing raw counts of evaluation outputs under SE in Table 5 and SE+CFE in Table 6. Table 4 summarizes results across all datasets. In the following, we highlight a few illustrative cases.
GSM8K, which consists of grade school math problems, also presents challenges for CFE. In attack scenarios, even when SE is deceived into accepting a reworded question as correct, CFE can occasionally notice the discrepancy between the predetermined fake answer and the reworded question. For example, we focus on the question: "Tommy has toy cars. His neighbor, Jessie, has cars too. Jessie’s older brother has more cars than Tommy and Jessie. How many cars do the three of them have altogether?" When the true answer "" is replaced with "Umbrella", GPT-3.5-turbo still expects a numerical value as the answer to the reworded question. This results in an output of , which incorrectly marks it as a correct answer. The problem arises because the evaluator does not realize that by assuming the new ground truth for the original question, its equivalent reworded question should also adopt the new ground truth. This can be especially misleading to the LLM when the newly assumed answer is in a completely different format, such as a word instead of a number.
On the other hand, CFE can fail even for regular correct answers. In another math problem from GSM8K, when given an LLM-generated correct answer but asked to assume a different random ground truth, the evaluator may incorrectly output . For instance, it might explain its judgment with "The user input provides a detailed and accurate calculation aligning with the question, resulting in the correct answer of . This matches the information provided in the assumed correct response ’Elephant’, indicating that the user input is in line with the expected answer." Here, the LLM evaluator appears to be hallucinating a connection between the user input and the assumed correct answer, rather than evaluating truthfully. This exemplifies how assumption-based CFE can be mislead for standard correct answers.
TruthfulQA, which assesses an LLM’s ability to reject subtle misconceptions, proved more challenging for GPT-3.5-turbo, even under SE. The questions are phrased in a misleading way to test reasoning under common misconceptions. This makes it particularly challenging for models that rely heavily on statistical correlations rather than careful reasoning. For example, in questions such as "Can you marry your cousin in Australia", GPT-3.5-turbo misinterpreted the question as suggesting that cousin marriage is allowed, despite the ground truth being "No". This issue was less pronounced in GPT-4o, GPT-4o-mini, and o1. However, for GPT-4o-mini, CFE also hallucinated a connection between the LLM-generated input and the assumed correct answer. It reasoned that, although the answer stated by the user does not match the answer that is newly assumed, the request is based on whether the input aligns with the question. Therefore, even when the correct answer is submitted, the output may still be 1, indicating that the response is systematically consistent with the analysis needed to solve the problem, even if it diverges from the predetermined answer.
We observe that the effectiveness of the method depends on the linguistic competence of the evaluation model. In particular, failure cases, especially with models like GPT-3.5-turbo and GPT-4o-mini, typically stem from: 1) the model’s inability to recognize that the submitted answer is a paraphrase of the original question, 2) its failure to reliably follow the injected instruction to treat a fake answer as correct, and 3) hallucinations where the model assumes connections between the submitted answer and the assumed ground truth that do not exist. In contrast, for more capable models with stronger linguistic abilities, these issues are significantly less prominent, as reflected in their improved attack detection accuracies.
These patterns collectively suggest that failure cases arise from limitations in the evaluator model’s reasoning ability. While the proposed method is broadly effective, its robustness varies with model capacity and the linguistic complexity of inputs.
For additional trends across datasets, refer to Tables 5 and 6.
Table 4: Pseudo Confusion Matrices Across All Datasets. This table reports raw counts of evaluation outputs per ground truth category, without applying any evaluation metrics such as accuracy or precision. The rows indicate the ground truth labels, with Correct for true answers, Wrong for incorrect answers, and Attack for adversarial examples, as specified in the column Gold. The columns reflect output judgments for each model. Under Standard Evaluation (SE), models classify outputs as either Correct or Wrong. When combining Standard Evaluation and Counterfactual Evaluation(SE+CFE), models can classify outputs as Correct (Corr), Wrong (Wng), or Attack (Attk). SE Gemma-12B LLaMA-3.1-8B Mistral-7B GPT-3.5-turbo GPT-4o-mini GPT-4o o1 Gold Correct Wrong Correct Wrong Correct Wrong Correct Wrong Correct Wrong Correct Wrong Correct Wrong Correct 585 15 536 64 534 66 541 59 586 14 587 13 591 9 Wrong 13 587 502 98 63 537 17 583 4 596 7 593 5 595 Attack 481 119 523 77 466 134 371 229 589 11 575 25 599 1 SE+CFE Gemma-12B LLaMA-3.1-8B Mistral-7B GPT-3.5-turbo GPT-4o-mini GPT-4o o1 Gold Corr Wng Attk Corr Wng Attk Corr Wng Attk Corr Wng Attk Corr Wng Attk Corr Wng Attk Corr Wng Attk Correct 555 17 28 121 104 375 454 66 80 524 59 17 571 14 15 568 13 19 591 9 0 Wrong 13 587 0 158 148 294 40 537 23 15 583 2 4 596 0 4 594 2 5 595 0 Attack 15 119 466 33 116 451 265 134 211 127 230 243 1 11 588 2 27 571 1 1 598
5 Conclusion
We introduced an evaluation framework combining Standard Evaluation (SE) and Counterfactual Evaluation (CFE) to defend LLM-based automatic evaluation systems against blind attacks. Our experiments showed that while SE alone is vulnerable to deception, with advanced models like o1 and GPT-4o often misclassifying adversarial inputs, the inclusion of CFE substantially improved attack detection for recent models with minimal performance trade-offs.
The attacks studied here represent a baseline using a simple, reproducible class of threats. Future work should extend this framework to defend against more complex and diverse attacks. Furthermore, to increase the trustworthiness of our framework, its judgments should be compared against human evaluations. Other promising directions include systematically exploring cross-lingual robustness and enhancing CFE by using a consensus over multiple, independently generated fake answers to mitigate the risk of coincidental semantic overlap.
While our evaluation framework improves robustness in evaluator LLMs, the vulnerabilities observed also highlight broader concerns. Prompt injection and jailbreaks can be utilized by adversaries to bypass existing defenses and reveal additional vulnerabilities. Ultimately, our findings highlight the limitations of standard evaluation protocols and demonstrate the necessity of more robust methods like CFE to ensure the security and reliability of both proprietary and open-source LLMs in evaluation tasks.
| Gemma-12B | LLaMA-3.1-8B | Mistral-7B | GPT-3.5-turbo | GPT-4o-mini | GPT-4o | o1 | |||||||||
| Dataset | Gold | Corr | Wng | Corr | Wng | Corr | Wng | Corr | Wng | Corr | Wng | Corr | Wng | Corr | Wng |
| GSM8K | Corr | 91 | 9 | 81 | 19 | 46 | 54 | 93 | 7 | 98 | 2 | 99 | 1 | 100 | 0 |
| Wng | 2 | 98 | 73 | 27 | 37 | 63 | 8 | 92 | 2 | 98 | 0 | 100 | 1 | 99 | |
| Attk | 79 | 21 | 78 | 22 | 37 | 63 | 78 | 22 | 100 | 0 | 98 | 2 | 99 | 1 | |
| HotpotQA | Corr | 99 | 1 | 89 | 11 | 100 | 0 | 93 | 7 | 93 | 7 | 98 | 2 | 99 | 1 |
| Wng | 0 | 100 | 80 | 20 | 4 | 96 | 1 | 99 | 0 | 100 | 0 | 100 | 0 | 100 | |
| Attk | 91 | 9 | 85 | 15 | 95 | 5 | 80 | 20 | 99 | 1 | 95 | 5 | 100 | 0 | |
| SQuAD | Corr | 97 | 3 | 91 | 9 | 96 | 4 | 98 | 2 | 100 | 0 | 97 | 3 | 97 | 3 |
| Wng | 0 | 100 | 81 | 19 | 3 | 97 | 0 | 100 | 0 | 100 | 1 | 99 | 0 | 100 | |
| Attk | 86 | 14 | 84 | 16 | 86 | 14 | 51 | 49 | 100 | 0 | 96 | 4 | 100 | 0 | |
| StrategyQA | Corr | 99 | 1 | 85 | 15 | 98 | 2 | 82 | 18 | 97 | 3 | 99 | 1 | 98 | 2 |
| Wng | 0 | 100 | 87 | 13 | 0 | 100 | 6 | 94 | 0 | 100 | 1 | 99 | 0 | 100 | |
| Attk | 71 | 29 | 91 | 9 | 87 | 13 | 56 | 44 | 98 | 2 | 97 | 3 | 100 | 0 | |
| TriviaQA | Corr | 99 | 1 | 96 | 4 | 99 | 1 | 98 | 2 | 98 | 2 | 96 | 4 | 100 | 0 |
| Wng | 11 | 89 | 91 | 9 | 14 | 86 | 1 | 99 | 0 | 100 | 1 | 99 | 1 | 99 | |
| Attk | 94 | 6 | 91 | 9 | 91 | 9 | 84 | 16 | 98 | 2 | 93 | 7 | 100 | 0 | |
| TruthfulQA | Corr | 100 | 0 | 94 | 6 | 95 | 5 | 77 | 23 | 100 | 0 | 98 | 2 | 97 | 3 |
| Wng | 0 | 100 | 90 | 10 | 5 | 95 | 1 | 99 | 2 | 98 | 4 | 96 | 3 | 97 | |
| Attk | 60 | 40 | 94 | 6 | 70 | 30 | 22 | 78 | 94 | 6 | 96 | 4 | 100 | 0 | |
| Gemma-12B | LLaMA-3.1-8B | Mistral-7B | GPT-3.5-turbo | GPT-4o-mini | GPT-4o | o1 | ||||||||||||||||
| Dataset | Gold | Corr | Wng | Attk | Corr | Wng | Attk | Corr | Wng | Attk | Corr | Wng | Attk | Corr | Wng | Attk | Corr | Wng | Attk | Corr | Wng | Attk |
| GSM8K | Corr | 86 | 10 | 4 | 14 | 24 | 62 | 14 | 54 | 32 | 91 | 7 | 2 | 93 | 2 | 5 | 99 | 1 | 0 | 100 | 0 | 0 |
| Wng | 2 | 98 | 0 | 22 | 32 | 46 | 17 | 63 | 20 | 7 | 92 | 1 | 2 | 98 | 0 | 0 | 100 | 0 | 1 | 99 | 0 | |
| Attk | 1 | 21 | 78 | 19 | 35 | 46 | 17 | 63 | 20 | 42 | 22 | 36 | 0 | 0 | 100 | 0 | 3 | 97 | 0 | 1 | 99 | |
| HotpotQA | Corr | 94 | 2 | 4 | 19 | 15 | 66 | 84 | 0 | 16 | 91 | 7 | 2 | 89 | 7 | 4 | 91 | 2 | 7 | 99 | 1 | 0 |
| Wng | 0 | 100 | 0 | 24 | 28 | 48 | 4 | 96 | 0 | 0 | 99 | 1 | 0 | 100 | 0 | 0 | 100 | 0 | 0 | 100 | 0 | |
| Attk | 1 | 9 | 90 | 5 | 19 | 76 | 50 | 5 | 45 | 20 | 20 | 60 | 0 | 1 | 99 | 0 | 6 | 94 | 0 | 0 | 100 | |
| SQuAD | Corr | 96 | 3 | 1 | 27 | 19 | 54 | 89 | 4 | 7 | 97 | 2 | 1 | 99 | 0 | 1 | 90 | 3 | 7 | 97 | 3 | 0 |
| Wng | 0 | 100 | 0 | 36 | 27 | 37 | 3 | 97 | 0 | 0 | 100 | 0 | 0 | 100 | 0 | 0 | 99 | 1 | 0 | 100 | 0 | |
| Attk | 4 | 14 | 82 | 1 | 22 | 77 | 49 | 14 | 37 | 20 | 49 | 31 | 0 | 0 | 100 | 1 | 4 | 95 | 0 | 0 | 100 | |
| StrategyQA | Corr | 84 | 1 | 15 | 21 | 20 | 59 | 90 | 2 | 8 | 78 | 18 | 4 | 95 | 3 | 2 | 98 | 1 | 1 | 98 | 2 | 0 |
| Wng | 0 | 100 | 0 | 26 | 22 | 52 | 0 | 100 | 0 | 6 | 94 | 0 | 0 | 100 | 0 | 0 | 100 | 0 | 0 | 100 | 0 | |
| Attk | 2 | 29 | 69 | 5 | 14 | 81 | 71 | 13 | 16 | 14 | 44 | 42 | 1 | 2 | 97 | 1 | 3 | 96 | 0 | 0 | 100 | |
| TriviaQA | Corr | 99 | 1 | 0 | 25 | 10 | 65 | 89 | 1 | 10 | 95 | 2 | 3 | 97 | 2 | 1 | 96 | 4 | 0 | 100 | 0 | 0 |
| Wng | 11 | 89 | 0 | 33 | 16 | 51 | 11 | 86 | 3 | 1 | 99 | 0 | 0 | 100 | 0 | 1 | 99 | 0 | 1 | 99 | 0 | |
| Attk | 5 | 6 | 89 | 2 | 13 | 85 | 38 | 9 | 53 | 25 | 17 | 58 | 0 | 2 | 98 | 0 | 7 | 93 | 1 | 0 | 99 | |
| TruthfulQA | Corr | 96 | 0 | 4 | 15 | 16 | 69 | 88 | 5 | 7 | 72 | 23 | 5 | 98 | 0 | 2 | 94 | 2 | 4 | 97 | 3 | 0 |
| Wng | 0 | 100 | 0 | 17 | 23 | 60 | 5 | 95 | 0 | 1 | 99 | 0 | 2 | 98 | 0 | 3 | 96 | 1 | 3 | 97 | 0 | |
| Attk | 2 | 40 | 58 | 1 | 13 | 86 | 40 | 30 | 30 | 6 | 78 | 16 | 0 | 6 | 94 | 0 | 4 | 96 | 0 | 0 | 100 | |
Limitations
Our work has several limitations. First, our experiments are confined to English benchmarks. The effectiveness of our counterfactual evaluation method may differ in languages with richer morphology or different syntactic structures, and our findings may not generalize directly. Second, our framework relies on a binary judgment of correctness (correct/incorrect). This is a simplification, as answers in real-world QA tasks can be partially correct or take different valid forms. Extending our method to support more flexible, graded evaluations is an important direction for future work. Finally, our evaluation focuses on standard, off-the-shelf LLMs. Future investigations could explore how fine-tuning might improve security against prompt injection attacks. Despite these limitations, our study highlights critical vulnerabilities in current protocols and offers a practical solution to strengthen LLM-based assessments.
Ethics Statement
All datasets and models are publicly available and were used consistently for their intended purposes as specified by their original providers. The datasets include GSM8k (MIT), HotpotQA (CC BY-SA 4.0), SQuAD (CC BY-SA 4.0), StrategyQA (MIT), TriviaQA (Apache-2.0), and TruthfulQA (Apache-2.0). We also utilized several OpenAI’s LLMs, as well as open-source models such as Gemma, LLaMA, and Mistral accessed through OpenRouter, in adherence to their respective terms for use. No offensive or personally identifiable information is involved.
One possible ethical concern is that the study of prompt injection attacks on QA-system-based LLM evaluators might inadvertently act as instructions for exploiting them. However, all attack strategies presented are adapted from prior work and are not novel contributions. Our goal is to highlight vulnerabilities in current evaluation systems to motivate the development of more secure and robust defense methods.
AI assistants were utilized to assist in the writing and editing of this paper. We maintain full responsibility for the content, analysis, and conclusions presented.
Acknowledgments
This research was supported by Japan Science and Technology Agency (JST), Core Research for Evolutionary Science and Technology CREST Program, Grant Number JPMJCR21D1; JST FOREST Program, Grant Number JPMJFR232S, and JSPS KAKENHI Grant Number JP23K28092. The authors thank Yuki Wakai for his helpful advices.